5%=96.21
44.15%=1443.88
64.17%=3270.09
100%=?

Answer by Cromlix(4381)

5%=96.21
44.15%=1443.88
64.17%=3270.09
100%=?
........
5% = 96.21
1% = 96.21/5
1% = 19.242
100% = 19.242 * 100
100% = 1924.2
........
44.15% = 1443.88
1% = 1443.88/44.15
= 32.704
100% = 32.704 * 100
100% = 3270.4
........
64.17% = 3270.09
1% = 50.960
100% = 50.960 * 100
100% = 5096
..................
Hope this helps.
:-)